Please help me answer this question.
Which of the following equations describes a line that passes through the points (2, 1) and (0, -5)?
A) Y= (1/3)x-5
B) Y= 3x-5
C) Y= -(1/3)x-5
D) Y= -3x-5

Answer by macston(5194)

.
ANSWER: B).
.
y=3x-5
.
Substitute values for point A (2,1)
x=2, y=1
y=3x-5
1=3(2)-5
1=6-5
1=1
.
Substitute values for Point B (0.-5)
x=0, y=-5
y=3x-5
-5=3(0)-5
-5=-5
.
You should try the other three yourself to make sure they won't work.
.